### v3

# Moving Mesh with Streamline Upwind Petrov-Galerkin (MM-SUPG) Method for Time-dependent Convection-Dominated Convection-Diffusion Problems

Abstract
Time-dependent convection-diffusion problems is considered, particularly when the diffusivity is very small and sharp layers exist in the solutions. Nonphysical oscillations may occur in the numerical solutions when using regular mesh with standard computational methods. In this work, we develop a moving mesh SUPG (MM-SUPG) method, which integrates the streamline upwind Petrov-Galerkin (SUPG) method with the moving mesh partial differential equation (MMPDE) approach. The proposed method is designed to handle both isotropic and anisotropic diffusivity tensors. For the isotropic case, we focus on improving the stability of the numerical solution by utilizing both artificial diffusion from SUPG and mesh adaptation from MMPDE. And for the anisotropic case, we focus on the positivity of the numerical solution. We introduce a weighted diffusion tensor and develop a new metric tensor to control the mesh movement. We also develop conditions for time step size so that the numerical solution satisfies the discrete maximum principle (DMP). Numerical results demonstrate that the proposed MM-SUPG method provides results better than SUPG with fixed mesh or moving mesh without SUPG.
1 Introduction
We consider the time-dependent convection-diffusion equation defined over the time-independent domain $\Omega\subset\mathbb{R}^{2}$ with a Lipschitz boundary and denote $\Omega_{T}=\Omega\times[0,T]$. The equation is given by:
[…]
Here, $\mathbb{D}=\mathbb{D}({\boldsymbol{x}})$ on $\Omega_{T}=\Omega\times(0,T]$ is the diffusivity tensor that is symmetric and positive-definite (SPD), ${\boldsymbol{b}}$ represents an incompressible velocity field, while $g$, $u_{0}$, and $f$ are sufficiently smooth functions. The parameter $\varepsilon\in(0,1]$ is the scaling factor for the diffusivity coefficient. When $\mathbb{D}$ has the same eigenvalues, the problem is isotropic. In the anisotropic case, the eigenvalues of $\mathbb{D}$ are different, which indicates that diffusion in one direction is faster than in the other.
Our focus lies in the convection-dominated scenario where $\varepsilon$ is much smaller than $\|{\boldsymbol{b}}\|_{L^{\infty}}$. Such problems arise in various fields, including pollutant dispersal in rivers, atmospheric pollution, the Stefan problem, turbulent transport, and are particularly intriguing in the investigation of the Navier–Stokes equations with large Reynolds numbers, see [1] for relevant applications. When $\varepsilon$ is small, non-physical oscillations tend to arise in the solution unless an extremely fine spatial-temporal discretization is employed, incurring substantial computational costs. To enhance stability, artificial diffusion in the direction of ${\boldsymbol{b}}$ is proposed in [2] to be added to the standard Galerkin finite element method.
The streamline upwind Petrov-Galerkin (SUPG) method has proven to be particular effective in numerically addressing boundary and interior layer problems in both time-dependent and time-independent scenarios [3, 4, 5, 6]. The motivation behind SUPG is to employ upwinding of the test functions along the streamlines, aiming to enhance the diffusion of solution. This concept is most clearly illustrated in the 1D time-independent case using finite difference stencils:
$\frac{\partial u}{\partial x}=-\varepsilon\frac{\partial^{2}u}{\partial x^{2}}.$ (2)
When applying the finite difference method with only the center stencil, one obtains a numerical approximation that, depending on $\varepsilon$, tends to be under-diffusive. Conversely, utilizing only an upwind stencil leads to excessive diffusion, known as over-diffusive. As the exact solution typically lies somewhere in between these extremes, it prompts the adoption of upwinding the center scheme [7]. This method was initially proposed by Brooks and Hughes for incompressible Navier-Stokes problems [2], and has since proven to be fruitful in studying those equations [8, 9]. Comparison among different methods for the steady-state form of equation (1) have been performed in [10].
Efforts have been made to numerically solve equation (1a) for time-independent solutions, particularly focusing on steady flows [3, 11], by employing adaptive mesh methods in conjunction with SUPG, especially when $\varepsilon\ll 1$. Notably, mesh adaptation has emerged as a potent tool in addressing convection-dominated problems. Our investigation targets a specific class of problems characterized by two key properties: artificial (non-physical) oscillations and sharp layers, encompassing both interior and boundary layers. Consequently, our interest lies in devising a mesh strategy grounded in anisotropic mesh adaptation techniques.
Several notable methodologies have been explored for solving equation (1). For instance, in [12], Lagrange-Galerkin error analysis demonstrated first-order uniform convergence in $\varepsilon$. Additionally, $\varepsilon$-uniform convergence theory for finite element method with homogeneous boundary conditions was investigated in [13]. Carey et al. introduced adaptive mesh refinement (AMR) in [14], while Chinchapatnam et al. explored unsymmetric and symmetric meshless collocation techniques with radial basis functions in [15]. Of particular interest are the so-called SOLD (Spurious Oscillations at Layers Diminishing) schemes, which have been extensively studied in [16, 17, 18]. These schemes involve the addition of either isotropic or anisotropic artificial diffusion to the Galerkin formulation. Furthermore, in [19], local discontinuous Galerkin (LDG) methods were employed on a class of layer-adapted Shishkin and Bakhvalov-type meshes, accompanied by error analysis, to approximate equation (1).
For time-dependent problems, Wang et al. utilized the singular boundary method with $f=0$ in [20], employing an exponential variable transform along with the Fourier transform. Additionally, unfitted finite element methods have gained traction for tackling equation (1) in time-dependent domains [21]. These types of unfitted FEMs are known by various names such as XFEM, CutFEM, Finite Cell Method, and TraceFEM (refer to the sources in [21] along with [22, 23]). It is worth noting, as mentioned in [24], that unfitted techniques may encounter conditioning difficulties, which are addressed using specialized finite element spaces and aggregation technique.
In all the above mentioned methods, only isotropic diffusivity tensors were considered. However, in many practical applications, the diffusivity tensor is anisotropic. When the mesh is not aligned in the principal diffusion direction, non-physical oscillations may occur when using standard computational methods [25]. Special care must be taken to ensure that the numerical solution satisfies the discrete maximum principle (DMP). For problem with only diffusion, metric tensors has been developed in [25] for time-independent problems to adapt the mesh such as $\mathbb{M}_{DMP}=\theta_{K}\mathbb{D}_{K}^{-1}$ based on DMP satisfaction and $\mathbb{M}_{DMP+adap}=C_{K}(det(\mathbb{D}))^{\frac{1}{d}}\mathbb{D}_{K}^{-1}$ based on DMP satisfaction as well as minimization of interpolation errors. For time-dependent problems, condition for time step size is also developed in [26] to ensure that the numerical solutions satisfy DMP. Lu et al. investigated the satisfaction of DMP for anisotropic diffusion-convection-advection problems in [27]. However, it is even more challenge for numerical solutions of convection-dominated convection-diffusion problems to satisfy the DMP, which will be addressed in this paper.
In this paper, we propose to integrate the moving mesh partial differential equation (MMPDE) method with streamline upwind Petro-Galerkin (SUPG) method and call it as moving mesh SUPG (MM-SUPG) method. It works for both time-independent and time-dependent flows characterized by initial sharp layers, and can deal with both isotropic and anisotropic diffusivity tensors. It has been observed that, even for isotropic diffusivity, oscillations can persist with adapted meshes [28]. The challenge in numerically solving equation (1) becomes evident particularly when $\varepsilon\ll 1$ and the initial profile experiences convection over time or when the profile features sharp regular, parabolic, corner, or interior layers that develop or exist initially. Meanwhile, when $\mathbb{D}$ is anisotropic, non-physical solutions may appear in the numerical solution, which is a violation of DMP [25, 26]. To address this issue, we introduce a metric tensor to control the mesh movement and develop conditions for mesh and time step size so that the numerical solution satisfies DMP.
The remainder of this paper is organized as follows. In Section 2, we give a brief introduction for SUPG and MMPDE, and propose our MM-SUPG method. Section 3 presents numerical examples for isotropic diffusivity case to demonstrate the effectiveness of our approach, while also delineating the constraints of the moving mesh method. In Section 4, we investigate the satisfaction of discrete maximum principle for anisotropic diffusion. Finally, a summary and some conclusions are presented in Section 5.
2 Moving Mesh with Streamline Upwind Petrov-Galerkin (MM-SUPG) Method
Firstly, we give a brief introduction to SUPG method and MMPDE method. Then we explain in details our MM-SUPG method, the integration of MMPDE with SUPG.
2.1 SUPG method
Consider the convection-diffusion equation (1a) with homogeneous Dirichlet boundary conditions (1b) and initial condition (1c). The SUPG method is a stabilized finite element method that introduces artificial diffusion to the Galerkin formulation. Let $L^{2}(\Omega)$ and $H^{k}(\Omega)$, $k\geq 1$ denote the standard Lebesgue and Sobolev spaces, respectively. Let $H_{g}^{1}(\Omega)=\{v\in H^{1}(\Omega),\,v|_{\partial\Omega}=g\}$ and $H_{0}^{1}(\Omega)=\{v\in H^{1}(\Omega),\,v|_{\partial\Omega}=0\}$. The weak formulation of (1a) is to find $u\in H_{g}^{1}(\Omega)$ such that
$\int_{\Omega}u_{t}v-\varepsilon\int_{\Omega}\nabla\cdot(\mathbb{D}\nabla u)v+\int_{\Omega}({\boldsymbol{b}}\cdot\nabla u)v=\int_{\Omega}fv,\quad\forall v\in H_{0}^{1}(\Omega).$ (3)
We consider an affine family of simplicial triangulations $\{\mathcal{T}_{h}\}$ for $\Omega$ and denote the piecewise linear finite element space as $V^{h}=(v\in H^{1}(\Omega),\,v|_{K}\in\mathcal{P}_{1}^{K},\forall K\in\mathcal{T}_{h})$, where $\mathcal{P}_{1}^{K}$ is the linear polynomial space in element $K$. Denote by $U^{h}=V^{h}\cap H_{g}^{1}(\Omega)$ and $U_{0}^{h}=V_{h}\cap H_{0}^{1}(\Omega)$. The standard Galerkin method is to find $u_{h}\in U^{h}$ such that
$\int_{\Omega}u_{t}v_{h}-\varepsilon\int_{\Omega}\nabla\cdot(\mathbb{D}\nabla u_{h})v_{h}+\sum_{K\in\mathcal{T}_{h}}({\boldsymbol{b}}\cdot\nabla u_{h})v_{h}=\int_{\Omega}fv_{h},\quad\forall v_{h}\in U_{0}^{h}.$ (4)
The essential idea of SUPG involves altering the test function space to operate within the realm of functions that upwind the trial functions along the streamlines. This can be represented as the space:
$U_{h}^{\textrm{SUPG}}=\left\{v_{h}+\tau_{K}{\boldsymbol{b}}\cdot\nabla v_{h}\mid v_{h}\in U_{0}^{h}\right\},$ (5)
where $\tau_{K}$ is the stabilization parameter. Therefore, one may perceive the space $U_{h}^{\textrm{SUPG}}$ as a suitable stabilized test space for singularly perturbed problems. The weak formulation for SUPG method is given as follows: find $u\in H^{1}(\Omega)$ such that
$\int_{\Omega}u_{t}v_{h}-\varepsilon\int_{\Omega}\nabla\cdot(\mathbb{D}\nabla u_{h})(v_{h}+\tau_{K}{\boldsymbol{b}}\cdot\nabla v_{h})+\sum_{K\in\mathcal{T}_{h}}({\boldsymbol{b}}\cdot\nabla u_{h})(v_{h}+\tau_{K}{\boldsymbol{b}}\cdot\nabla v_{h})=\int_{\Omega}f(v_{h}+\tau_{K}{\boldsymbol{b}}\cdot\nabla v_{h}),\quad\forall v_{h}\in U_{0}^{h}.$ (6)
A conventional choice of $\tau_{K}$ on mesh $\mathcal{T}_{h}$ is
$\tau_{\mathcal{T}_{h}}^{K}=\begin{cases}c_{0}\textrm{diam}(K)_{\mathcal{T}_{h}}\ \ \textrm{if}\ \ \textrm{Pe}_{K}>1\text{ convection-dominated}\\ c_{1}\textrm{diam}(K)^{2}_{\mathcal{T}_{h}}\ \ \textrm{if}\ \ \textrm{Pe}_{K}\leq 1\text{ diffusion-dominated}\end{cases},$ (7)
where $\textrm{Pe}_{K}$ denotes the element-wise Péclet number given by
$\textrm{Pe}_{K}=\frac{\|{\boldsymbol{b}}\|_{\infty}\textrm{diam}(K)_{\mathcal{T}_{h}}}{2\varepsilon}.$ (8)
It determines local convection-dominated or diffusion-dominated regimes on the current mesh $\mathcal{T}_{K}$. There is flexibility in defining $\textrm{diam}(K)_{\mathcal{T}_{h}}$ and in selecting $c_{0}$ and $c_{1}$. This paper primarily focuses on scenarios where $\textrm{Pe}_{K}$ is large with respect to $\varepsilon^{-1}\|{\boldsymbol{b}}\|_{\infty}$. Typically, individual bilinear forms in the SUPG formulation are estimated, and $\tau_{\mathcal{T}_{h}}^{K}$ is derived based on those errors. In [3], such a method was employed using error estimates in the discrete norm
$\|w\|_{\varepsilon,k}^{2}=\varepsilon\|\nabla w\|^{2}+\sum_{K\in\mathcal{T}_{K}}\tau_{\mathcal{T}_{h}}^{K}\|{\boldsymbol{b}}\cdot\nabla w\|_{L^{2}(K)}^{2},\,\forall w\in H_{0}^{1}(\Omega)$ (9)
for time independent ${\boldsymbol{b}}$. However, our method relies on different error estimates for defining the metric tensor along with a mesh functional approach. If a source term is introduced, say $c\in L^{\infty}(0,T;L^{\infty}(\Omega))$, it has been proposed that the FE perturbation parameters should consider this source term [29, 30, 31, 32].
In the case of anisotropic mesh adaptation, $\textrm{diam}(K)_{\mathcal{T}_{h}}$ will be updated with each new mesh as the elements stretch and shrink. In this paper, we define $\tau_{\mathcal{T}_{h}}^{K}$ as
$\tau_{\mathcal{T}_{h}}^{K}$ $=\frac{\textrm{diam}(K)_{\mathcal{T}_{K}}}{2\|{\boldsymbol{b}}\|_{\infty}}\xi(\textrm{Pe}_{K}),$ (10)
which is based on the stabilization parameter selected in [11]. Common choices of $\textrm{diam}(K)_{\mathcal{T}_{h}}$ include the length of the longest edge of the element $K$ projected onto ${\boldsymbol{b}}$ and the diameter of $K$ in the direction of ${\boldsymbol{b}}$. In this paper, for each $K\in\mathcal{T}_{h}$, we define
$\textrm{diam}(K)=\sup_{{\boldsymbol{x}}_{1},{\boldsymbol{x}}_{2}\in K}\|{\boldsymbol{x}}_{1}-{\boldsymbol{x}}_{2}\|$ (11)
and
$\xi(\textrm{Pe}_{K})=\min\left\{1,\frac{\textrm{Pe}_{K}}{3}\right\}.$ (12)
2.2 Moving mesh partial differential equations (MMPDE)
Moving mesh methods are instrumental in addressing domains where the solution exhibits excessively large gradient, such as sharp layers. Traditional methods like fitted operator methods, finite difference methods, finite element methods with uniform or Shishkin-type meshes often struggle to fully resolving these issues without a hyper-refined mesh or a priori knowledge about the solution. The MMPDE approach automatically concentrates mesh elements near the region where the solution exhibits large gradient, hence improves the computational efficiency. In this subsection, we give a brief overview of MMPDE. More details can be found in [33].
We consider an affine family of simplicial triangulations $\{\mathcal{T}_{h}\}$ for $\Omega$. Denote by $(K,\mathcal{P}_{1}^{K},\Sigma_{K})$ the finite element space. Let $N$, $N_{v}$, and $N_{t}$ denote the number of mesh elements, the number of vertices, and the number of time steps, respectively. Let $\mathbb{M}=\mathbb{M}({\boldsymbol{x}})$ be a symmetric and uniformly positive definite metric tensor defined on $\Omega$ with $c_{1}\mathbb{I}\leq\mathbb{M}({\boldsymbol{x}})\leq c_{2}\mathbb{I}$ for all ${\boldsymbol{x}}\in\Omega$. We take the $\mathbb{M}$-uniform mesh approach that takes a nonuniform mesh and views it as a uniform one in the metric specified by $\mathbb{M}$. Essentially, $\mathbb{M}$ controls the concentration and alignment of the mesh elements.
We consider two affine-equivalent finite elements, one for the reference elements and another for the physical mesh element. There exists an invertible affine mapping between the finite elements, denoted by $F_{K}$:
$(\hat{K},\mathcal{P}_{1}^{\hat{K}},\Sigma_{\hat{K}})\xrightarrow{\quad F_{K}\quad}(K,\mathcal{P}_{1}^{K},\Sigma_{K}),$ (13)
For each element $K\in\mathcal{T}_{h}$, denote by $F_{K}:\hat{K}\to K$ the affine mapping between $K$ and the reference element $\hat{K}$ so that $\lvert\hat{K}\rvert=1$. Denote by ${\boldsymbol{x}}_{j}^{K}$ for $j=0$, 1, and 2 the vertices of $K$ and ${\boldsymbol{\xi}}_{j}^{K}$ the corresponding vertices of $\hat{K}$. Then the affine map $F_{K}$ maps ${\boldsymbol{\xi}}_{j}^{K}$ to ${\boldsymbol{x}}_{j}^{K}$, i.e. $F_{K}({\boldsymbol{\xi}}_{j}^{K})={\boldsymbol{x}}_{j}^{K}$ for $j=0$, 1, and 2. $\hat{K}$ is called the reference element and belongs to $\hat{\mathcal{T}}_{h}$, the reference mesh.
Denote by $\mathbb{M}_{K}$ the volume-average of $\mathbb{M}$ over element $K$, i.e
$\mathbb{M}_{K}=\frac{1}{\lvert K\rvert}\int_{K}\mathbb{M}({\boldsymbol{x}})\textrm{d}{\boldsymbol{x}},$
and
$\sigma_{h}=\sum_{K\in\mathcal{T}_{h}}\int_{K}\sqrt{\det(\mathbb{M}_{K})}\textrm{d}{\boldsymbol{x}}=\sum_{K\in\mathcal{T}_{h}}\lvert K\rvert\sqrt{\det(\mathbb{M}_{K})}.$
$F_{K}$ and $\mathbb{M}$ give rise to the so-called equidistribution and alignment conditions in $\mathbb{R}^{2}$:
  1. 1.
    equidistribution: $\lvert K\rvert\sqrt{\det(\mathbb{M}_{K})}=\displaystyle\frac{\sigma_{h}}{N}$ for all $K\in\mathcal{T}_{h}$ and
  2. 2.
    alignment: $\frac{1}{2}\textrm{tr}\left(F_{K}^{\prime}\mathbb{M}_{K}(F_{K}^{\prime})^{T}\right)=\sqrt{{\textrm{det}}\left(F_{K}^{\prime}\mathbb{M}_{K}(F_{K}^{\prime})^{T}\right)}$ for all $K\in\mathcal{T}_{h}$,
where $F_{K}^{\prime}$ is the Jacobian matrix of $F_{K}$. These two completely characterize a non-uniform mesh: the equidistribution condition controls the size and shape of mesh elements, while the alignment condition controls the orientation of mesh elements. The new physical mesh $\mathcal{T}_{h}$ is found by minimizing the energy functional (also called the mesh functional) $I$ given in discrete form by
$I(\mathcal{T}_{h})=\sum_{K\in\mathcal{T}_{h}}\lvert K\rvert G(\mathbb{J}_{K},\det(\mathbb{J}_{K}),\mathbb{M}_{K}),$ (14)
where $\mathbb{J}_{K}=(F_{K}^{\prime})^{-1}$ and
$G(\mathbb{J}_{K},\det(\mathbb{J}_{K}),\mathbb{M}_{K})=$ $\alpha\sqrt{\det(\mathbb{M}_{K})}\left(\textrm{tr}\left(\mathbb{J}_{K}\mathbb{M}_{K}\mathbb{J}_{K}^{T}\right)\right)^{p}$ (15)
$+(1-2\alpha)2^{p}\sqrt{\det(\mathbb{M}_{K})}\left(\frac{\det(\mathbb{J}_{K})}{\sqrt{\det(\mathbb{M}_{K})}}\right)^{p},$
with $\alpha\in(0,\frac{1}{2}]$ and $p>1$ as dimensionaless parameters. From numerical experiments, $\alpha=\frac{1}{3}$ and $p=\frac{3}{2}$ work well for most problems [34].
The moving mesh PDE (MMPDE) method is developed by Huang et al. [33] to find a new physical mesh by minimizing $I(\mathcal{T}_{h})$, which is equivalent to the steepest descent method [35]. The MMPDE is a partial differential equation (also called the mesh PDE) that is given by
$\frac{\textrm{d}{\boldsymbol{x}}_{i}}{\textrm{d}t}=-\frac{P_{j}}{\gamma}\left(\frac{\partial I}{\partial{\boldsymbol{x}}_{i}}\right)^{T},\ i=1,\dots,N_{v},$ (16)
where $\gamma$ is a user parameter meant to adjust the time scale of the mesh movement and $P_{j}=\det\left(\mathbb{M}(x_{j})\right)^{\frac{p-1}{2}}$ is chosen so that the MMPDE is invariant under the scaling transformation of $\mathbb{M}$. The physical PDE and mesh PDE can be solved either simultaneously or alternatively. When the PDEs are solved simultaneously, the mesh can respond quickly to any change in the physical solution. However, this procedure is highly nonlinear. In our work, we employ the alternative approach. The mesh PDE is firstly solved based on the physical solution and the mesh at the current time step. Then mesh nodes are relocated based on the solutions from mesh PDE. Lastly, the physical PDE is solved on the new mesh for the next time step. The alternate procedure allows for flexibility and potential efficiency at each time level. One disadvantage of this approach is a lag in time, which can be addressed by employing multiple iterations on the same time level before advancing to the next. It is always beneficial to start with a initially adapted mesh, especially when sharp layers exist from the beginning.
In this paper, for moving mesh for isotropic diffusion case, we use the metric tensor defined by
$\mathbb{M}=\det\left(\mathbb{I}+\lvert H_{K}(u)\rvert\right)^{-\frac{1}{6}}\left(\mathbb{I}+\lvert H_{K}(u)\rvert\right),$ (17)
where $\mathbb{I}$ is the identity matrix, $H_{K}(u)$ is the recovered Hessian using least squares fitting to the values of the solution function at the mesh vertices, and $\lvert H_{K}(u)\rvert=\Lambda\textrm{diag}(\lvert\lambda_{1}\rvert,\lvert\lambda_{2}\rvert)\Lambda^{T}$ with $\Lambda\textrm{diag}(\lambda_{1},\lambda_{2})\Lambda^{T}$ being the eigen-decomposition of $H_{K}(u)$. The Hessian is computed $H(u)({\boldsymbol{x}}_{j})\approx H(q)({\boldsymbol{x}}_{j})$, where $q$ is the algebraic polynomial determined by least squares fitting. For completeness, some details for the least squares fitting method (see Section 5.3 in [33]) are provided here.
Specifically, if $u_{j}$ is the approximation of $u$ at node ${\boldsymbol{x}}_{j}=(x_{j},y_{j})$ on mesh $\mathcal{T}_{h}$. Let ${\boldsymbol{x}}_{j_{i}}$ for $i=1,\dots,N_{j}$ be the $N_{j}$ neighboring vertices near ${\boldsymbol{x}}_{j}$ (including ${\boldsymbol{x}}_{j}$). Denote by ${\boldsymbol{x}}_{j}^{c}$ the center of these points and define
$H_{j}^{x}=\max_{i=1,\dots,N_{j}}\lvert x_{j_{i}}-x_{j}^{c}\rvert,\ \ H_{j}^{y}=\max_{i=1,\dots,N_{j}}\lvert y_{j_{i}}-y_{j}^{c}\rvert.$ (18)
Then the polynomial $q$ in (Legendre) polynomial form
$q(x,y)=\sum_{k=0}^{2}\sum_{m=0}^{2-k}a_{k,m}P_{k}\left(\frac{x-x_{j}^{c}}{H_{j}^{x}}\right)P_{m}\left(\frac{y-y_{j}^{c}}{H_{j}^{y}}\right),$ (19)
and is determined by
$\min_{a_{0,0},\dots,,a_{2,0}}\sum_{i=1}^{N_{j}}(q(x_{j_{i}},y_{j_{i}})-u_{j_{i}})^{2}.$ (20)
Then $\nabla u({\boldsymbol{x}}_{j})\approx\nabla q({\boldsymbol{x}}_{j})$. This metric tensor is optimal in the $L^{2}$ norm of linear interpolation error on triangular meshes [35]. Moreover, it is clearly symmetric and one may replace $\lvert H_{K}(u)\rvert$ with $\alpha^{-1}|H_{K}(u)\rvert$, where $\alpha$ is called the intensity parameter and helps ensure that $\mathbb{M}$ is positive definite. This monitor function is derived on anisotropic error estimates (see [33], chapter 5, for a more detailed discussion and the derivation of (17) and choice of $\alpha$).
2.3 Moving mesh with Streamline Upwind Petrov-Galerkin (MM-SUPG) method
Now we apply the MMPDE method to solve the SUPG formulation (6). Write ${\boldsymbol{u}}_{h}({\boldsymbol{x}},t)=\sum\limits_{j=1}^{N}u_{j}(t)\varphi_{j}({\boldsymbol{x}})$, where $\{\varphi_{j}\}_{j=1}^{N}$ is the set of linear basis functions in $U_{h}^{0}$. Also, since $\varphi_{j}$ is a linear function, the element-wise integral average can be defined explicitly
$\frac{1}{|K|}\int_{K}\varphi_{j}\varphi_{i}\textrm{d}{\boldsymbol{x}}=\begin{cases}\displaystyle\frac{1}{(d+1)(d+2)}\textrm{\ \ for \ \ }i\neq j\\ \displaystyle\frac{2}{(d+1)(d+2)}\textrm{\ \ for \ \ }i=j\end{cases}.$ (21)
Approximate the global integral as
$\int_{\Omega}g({\boldsymbol{x}},t)\textrm{d}{\boldsymbol{x}}$ $=\sum_{K\in\mathcal{T}_{h}}|K|\int_{K}g({\boldsymbol{x}},t)\textrm{d}{\boldsymbol{x}}$ (22)
$\approx\sum_{K\in\mathcal{T}_{h}}|K|\sum_{j=1}^{m}\lambda_{j}g({\boldsymbol{\omega}}_{j}),$ (23)
where ${\boldsymbol{\omega}_{j}}$ are the quadrature nodes for element $K$ and $\lambda_{j}$ are weights such that $\sum_{j=1}^{m}\lambda_{j}=1$. For any tensor $\mathbb{F}$, we then define the element-wise average by
$\mathbb{F}_{K}=\frac{1}{|K|}\int_{K}\mathbb{F}({\boldsymbol{x}})\textrm{d}{\boldsymbol{x}}\approx\sum_{j=1}^{N}\lambda_{j}\mathbb{F}({\boldsymbol{\omega}_{j}}).$ (24)
The semi-discrete matrix formulation for (6) is given by
${\boldsymbol{M}}{\boldsymbol{u}}_{h}^{\prime}(t)+{\boldsymbol{A}}(t){\boldsymbol{u}}_{h}(t)={\boldsymbol{f}}(t),$ (25)
where ${\boldsymbol{u}}=[u_{1},\dots,u_{Nv}]^{T}$, ${\boldsymbol{M}}$, ${\boldsymbol{A}}$, and ${\boldsymbol{b}}$ are given by
$M_{ij}$ $=\sum_{K\in\mathcal{T}_{h}}(\varphi_{j},\varphi_{i})_{0,K}+\tau_{K}(\varphi_{j},{\boldsymbol{b}}\cdot\nabla\varphi_{i})_{0,K}$ (26)
$A_{ij}(t)$ $=\sum_{K\in\mathcal{T}_{h}}({\boldsymbol{b}}\cdot\nabla\varphi_{j},\varphi_{i})_{0,K}+|K|(\nabla\varphi_{i})^{T}\varepsilon\mathbb{D}_{K}\nabla\varphi_{j}$ (27)
$\quad\quad+\tau_{K}\left(({\boldsymbol{b}}\cdot\nabla\varphi_{j},{\boldsymbol{b}}\cdot\nabla\varphi_{i})_{0,K}+(\varepsilon\mathbb{D}_{K}\nabla\varphi_{j},\nabla({\boldsymbol{b}}\cdot\nabla\varphi_{i}))_{0,K}\right)$
$f_{i}(t)$ $=\sum_{K\in\mathcal{T}_{h}}(f,\varphi_{i})_{0,K}+\tau_{K}(f,{\boldsymbol{b}}\cdot\nabla\varphi_{i})_{0,K}.$ (28)
Denote $N_{t}$ as the number of time steps. We use a $\theta$ scheme to discretize in time. For (25), we approximate
${\boldsymbol{u}}_{h}^{\prime}(t)\approx\frac{{\boldsymbol{u}}_{h}(t_{m+1})-{\boldsymbol{u}}_{h}(t_{m})}{\Delta t},\ \ m=1,\dots,N_{t}.$ (29)
Then the corresponding $\theta$-scheme is given by
$M\frac{{\boldsymbol{u}}_{h}(t_{m+1})-{\boldsymbol{u}}_{h}(t_{m})}{\Delta t}$ $+\theta A{\boldsymbol{u}}_{h}(t_{m+1})+(1-\theta)A{\boldsymbol{u}}_{h}(t_{m})$ (30)
$=\theta{\boldsymbol{f}}(t_{m+1})+(1-\theta){\boldsymbol{f}}(t_{m}),\ \ m=1,\dots,N_{t}.$
Grouping the ${\boldsymbol{u}}_{h}(t_{m+1})$ terms, we may rewrite (30) as
$\left(\frac{M}{\Delta t}+\theta A\right){\boldsymbol{u}}_{h}(t_{m+1})=\theta{\boldsymbol{f}}(t_{m+1})+(1-\theta){\boldsymbol{f}}(t_{m})+\left(\frac{M}{\Delta t}-(1-\theta)A\right){\boldsymbol{u}}_{h}(t_{m}),$ (31)
for $m=1,\dots,N_{t}$. Hence, we may write a system of ODEs
$\tilde{{\boldsymbol{A}}}{\boldsymbol{u}}_{h}(t_{m+1})=\tilde{\boldsymbol{f}}^{m+1},\ \ m=1,\dots,N_{t},$ (32)
where
$\tilde{{\boldsymbol{A}}}$ $={\boldsymbol{M}}+\theta\Delta t{\boldsymbol{A}}$ (33)
$\tilde{\boldsymbol{f}}^{m+1}$ $=\theta\Delta t{\boldsymbol{f}}(t_{m+1})+\Delta t(1-\theta){\boldsymbol{f}}(t_{m})+\left({\boldsymbol{M}}-\Delta t(1-\theta){\boldsymbol{A}}\right){\boldsymbol{u}}_{h}(t_{m}).$ (34)
In this work, we use $\theta=0.5$, which corresponds to the Crank-Nicholson. Fixed time step is used in the computations. In Section 2 we establish bounds on $\Delta t$ so that DMP is satisfied.
3 Results for Isotropic Diffusion
Now, we present some numerical examples for isotropic diffusion case where we take the diffusivity tensor $\mathbb{D}$ as an identity matrix. The value of diffusivity is given by $\varepsilon$. We consider two different examples and compare results from four different methods: fixed mesh finite element method (FM-FEM) without SUPG, fixed mesh finite element method with SUPG (FM-SUPG), moving mesh finite element method (MM-FEM) without SUPG (which is the MMPDE method), and our moving mesh with SUPG (MM-SUPG) method. We denote the MMPDE method as MM-FEM for consistency in what follows.
      Example 3.1.
This example is similar to the simulated example presented in [36]. The domain is $\Omega=(0,10)\times(-1.25,1.25)$ and an initial smooth solution is given by
$u_{0}(x,y)=\textrm{e}^{-\frac{(x-1)^{2}+y^{2}}{0.25}}+\textrm{e}^{-\frac{(x-1)^{2}+(y-0.5)^{2}}{0.25}}+\textrm{e}^{-\frac{(x-1)^{2}+(y+0.5)^{2}}{0.25}},$ (35)
The velocity field is given by
${\boldsymbol{b}}=\left(1-\exp(\theta x)\cos(2\pi y),\displaystyle\frac{\theta}{2\pi}\exp(\theta x)\sin(2\pi y)\right),$ (36)
with
$\theta=\frac{1}{2\varepsilon}-\sqrt{\left(\frac{1}{2\epsilon}\right)^{2}+4\pi^{2}}.$ (37)
Refer to caption (a): FM-FEM           Refer to caption (b): FM-SUPG Refer to caption (c): MM-FEM           Refer to caption (d): MM-SUPG
Figure 1: Example 3.1 – side view of four different methods with $\Delta t=0.01$, ${\boldsymbol{b}}$ is given by (36), $N=32,768$ at $T=5.0$.
For this example, $\varepsilon=10^{-6}$ is used in the computation, which gives $\theta\approx-3.948\times 10^{-5}$. We compute the diffusion of the initial solution from 0 to 5 seconds with Dirichlet boundary conditions. Time step $\Delta t=10^{-2}$ and $N=32,768$ elements are used in the computations. Allowing the convection process to be carried out, we see fine layers form on the interior and boundary of $\Omega$.
The results are shown in Figure 1 with side view profiles and Figure 2 with top view profiles. From Figures 1(a) and (c), we see that for the classic Galerkin method, both FM-FEM and MM-FEM, the numerical solutions have non-physical oscillations form near the outflow boundary at $x=10$. Figure 1(b) shows that FM-SUPG resolves the boundary layers present in FM-FEM and MM-FEM but has artificial oscillations on the interior of $\Omega$. On the other hand, Figure 1(d) shows that MM-SUPG results in the smoothest surface among all four methods, and both interior oscillations and oscillations near the outflow have been resolved. Even with smaller $\Delta t$ as $5\times 10^{-3}$, we see in Figure 2 that MM-FEM still has layers at the outflow, where MM-SUPG performs much better.
Refer to caption (a): MM-FEM           Refer to caption (b): MM-SUPG
Figure 2: Example 3.1 – side view of MM-FEM and MM-SUPG methods with $\Delta t=0.005$, ${\boldsymbol{b}}$ is given by (36), $N=32,768$ at $T=5.0$.
      Example 3.2.
Here we consider another example (taken from [16]) with exact solution
$u(x,y,t)=1$ $6\sin(\pi t)x(1-x)y(1-y)$
×$\times\left(\frac{1}{2}+\frac{1}{\pi}\arctan\left(2\varepsilon^{-1/2}(0.25^{2}-(x-0.5)^{2}-(y-0.5)^{2})\right)\right),$ (38)
with Dirichlet boundary conditions. We compute to time $T=0.5$ with time step $\Delta t=10^{-3}$. The severity of the interior layer is proportional to $\varepsilon$ and its thickness is $\sqrt{\varepsilon}$. We also use ${\boldsymbol{b}}=(2,3)$ and $\varepsilon=10^{-6}$.
For coarse mesh with fewer elements (that is, smaller $N$), we observe artificial oscillations in the direction of ${\boldsymbol{b}}$, including MM-FEM, for example, with $N=8192$. Increasing the number of mesh elements helps to improve the results for MM-FEM. Figure 3 shows the top view of the solutions at $T=0.5$ for the four different methods with $N=32768$, while Figure 4 displays the side view of the solution profiles. As can be seen, significant smearing occurs in the solution obtained using the FM-FEM and FM-SUPG, though the latter shows an improvement. On the other hand, both the MM-FEM and MM-SUPG perform well with $N=32768$ elements in the mesh. This is supported by the $H^{1}$-seminorm errors in Table 1 and cap view in Figure 5. The $L^{2}$-norm errors in Table 2 shows that MM-SUPG performs a bit better than MM-FEM for this example. We observe pronounced mesh movement in the direction of ${\boldsymbol{b}}$ for MM-FEM and MM-SUPG, as shown in Figure 6. The rate of convergence in the $H^{1}$-seminorm and $L^{2}$ norm are shown in Figure 7.
Refer to caption (a): FM-FEM           Refer to caption (b): FM-SUPG Refer to caption (c): MM-FEM           Refer to caption (d): MM-SUPG
Figure 3: Example 3.2 – top view with four different methods with $\Delta t=0.001$ for ${\boldsymbol{b}}=(2,3)$, $N=32768$ at $T=0.5$.
Refer to caption (a): FM-FEM           Refer to caption (b): FM-SUPG Refer to caption (c): MM-FEM           Refer to caption (d): MM-SUPG
Figure 4: Example 3.2 – side view with the four different methods with $\Delta t=0.001$, ${\boldsymbol{b}}=(2,3)$, $N=32768$ at $T=0.5$.
Refer to caption (a): MM-FEM           Refer to caption (b): MM-SUPG
Figure 5: Example 3.2 – cap view moving mesh with $\Delta t=0.001$, ${\boldsymbol{b}}=(2,3)$, $N=32768$ at $T=0.5$.
Refer to caption (a): MM-FEM           Refer to caption (b): MM-SUPG
Figure 6: Example 3.2 – Mesh movement for $T=0.5$ and ${\boldsymbol{b}}=(2,3)$, $N=32768$ at $T=0.5$.
Refer to caption (a): $H^{1}$-seminorm error.           Refer to caption (b): $L^{2}$ norm error.
[…]
$N$ MM-FEM MM-SUPG FM-FEM FM-SUPG
512 12.1038 10.9715 27.6628 18.3147
2048 10.5624 11.0912 26.7879 15.3485
8192 4.4580 4.3027 34.9201 14.7499
32768 0.7387 0.9278 39.3491 10.7836
131072 0.2657 0.2101 39.2190 7.7920
[…]
$N$ MM-FEM MM-SUPG FM-FEM FM-SUPG
512 0.2633 0.1289 1.5475 0.4138
2048 0.1127 0.0823 0.8708 0.2871
8192 0.0475 0.0096 0.5818 0.1573
32768 0.0019 0.0010 0.3720 0.0510
131072 0.0004 0.0003 0.1995 0.0180
Within this example, we also consider a time-dependent flow given by ${\boldsymbol{b}}=(y-t,x-t)$. The convergence rates in the $H^{1}$-seminorm and $L^{2}$ norm can be found in Figure 8, where numerical values are recorded in Tables 3 and 4. Again, both MM-FEM and MM-SUPG perform much better than FM-FEM and FM-SUPG with larger $N$.
Refer to caption (a): $H^{1}$-seminorm error.           Refer to caption (b): $L^{2}$ norm error.
[…]
$N$ MM-FEM MM-SUPG FM-FEM FM-SUPG
512 11.1473 11.7017 16.7382 16.6045
2048 10.6529 9.7031 13.0265 12.8651
8192 5.8337 4.7036 12.9658 12.3947
32768 0.7614 0.9654 11.0075 9.8446
131072 0.1614 0.2054 8.7996 7.1405
[…]
$N$ MM-FEM MM-SUPG FM-FEM FM-SUPG
512 0.1629 0.1181 0.3340 0.1843
2048 0.0662 0.0530 0.2326 0.1193
8192 0.0392 0.0130 0.1724 0.0903
32768 0.0015 0.0011 0.1175 0.0411
131072 0.0002 0.0002 0.0602 0.0161
We see small oscillations along the top of the surface in the direction of the outflow in Figures 5 and 9, even for large $N$. However, for larger $N$, we see that DMP is not violated. Both methods converge to the same approximate solution as $N$ increases. In Figure 4 we see that the MM-SUPG method is a noticeable improvement compared to all methods. In Figure 9, only slight oscillations are observable on the backside of the surface (in the direction of the streamlines) for MM-FEM. This is more prominent for smaller $N$ in the MM-FEM case. Overall, MM-SUPG performs the best among all the four methods under consideration.
Refer to caption (a): MM-FEM           Refer to caption (b): MM-SUPG Refer to caption (c): Enlarged view of (a)           Refer to caption (d): Enlarged view of (b)
Figure 9: Example 3.2 – side view with $\Delta t=0.001$ and ${\boldsymbol{b}}=(y-t,x-t)$, $\varepsilon=10^{-6}$, and $N=32768$ for MM-FEM and MM-SUPG at $T=0.5$.
4 Discrete Maximum Principle for Anisotropic Diffusion
In this section, we consider the anisotropic diffusion case where the diffusivity tensor $\mathbb{D}$ has different eigenvalues, which indicates diffusion is faster in one direction (the principal diffusion direction) than in the other direction.
We say the solution $u$ to (1) satisfies the weak maximum principle if $u\in C^{2,1}(\Omega_{T})\cap C(\overline{\Omega}_{T})$ and $f\leq 0$ in $\Omega_{T}$ [37]:
$\max_{({\boldsymbol{x}},t)\in\overline{\Omega}_{T}}u\leq\max\left\{0,\max_{({\boldsymbol{x}},t)\in\Gamma_{T}}u({\boldsymbol{x}},t)\right\},$ (39)
where $\Gamma_{T}=\overline{\Omega}_{T}-\Omega_{T}$ is the parabolic boundary.
Let $S_{i}^{K}$ be the face opposite to the vertex ${\boldsymbol{a}}_{i}^{K}$, the $i$th vertex of element $K$. Introduce ${\boldsymbol{q}}$ vectors as [26]:
${\boldsymbol{q}}_{i}^{K}=\frac{{\boldsymbol{n}}_{i}^{K}}{h_{i}^{K}},$ (40)
where ${\boldsymbol{n}}_{i}^{K}$ is the unit inward normal pointing to ${\boldsymbol{a}}_{i}^{K}$, and $h_{i}^{K}$ is the distance from vertex ${\boldsymbol{a}}_{i}$ to surface $S_{i}$. Figure 10 shows an illustration of ${\boldsymbol{q}}$ vectors.
[…]
Figure 10: Illustration of faces and ${\boldsymbol{q}}$ vectors.
In [38] it was shown that
$\nabla\varphi_{i}|_{K}={\boldsymbol{q}}_{i}^{K}.$ (41)
Therefore,
$\int_{K}(\nabla\varphi_{i})^{2}\textrm{d}{\boldsymbol{x}}=\|{\boldsymbol{q}}\|^{2}|K|=\frac{|K|}{(h_{i}^{K})^{2}}.$ (42)
We also have the following inequality due to the Divergence Theorem:
$\int_{K}\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}=-\frac{1}{2}\int_{K}(\nabla\cdot{\boldsymbol{b}})\varphi_{i}^{2}\textrm{d}{\boldsymbol{x}}.$ (43)
Following [26], we define the diffusion-dependent map $\mathbb{D}_{K}^{-1/2}{\boldsymbol{x}}:K\to\tilde{K}$. Define the $\mathbb{D}_{K}$-norm by $\|{\boldsymbol{x}}\|_{\mathbb{D}_{K}}=\sqrt{{\boldsymbol{x}}^{T}\mathbb{D}_{K}{\boldsymbol{x}}}$. Then we have the following relations
$\tilde{\boldsymbol{a}}_{i}^{K}=\mathbb{D}_{K}^{-1/2}{\boldsymbol{a}}_{i}^{K},\ \ \tilde{S}_{i}^{K}=\mathbb{D}_{K}^{-1/2}S_{i}^{K},\ \ |\tilde{K}|=\det(\mathbb{D}_{K})^{1/2}|K|,\ \ \tilde{\boldsymbol{q}}_{i}^{K}=\mathbb{D}_{K}^{1/2}{\boldsymbol{q}}_{i}^{K},\ \ \tilde{h}_{i}^{K}=\|\tilde{\boldsymbol{q}}_{i}^{K}\|^{-1}.$ (44)
Denote by $\tilde{\alpha}_{ij}^{K}$ the dihedral angle between surfaces $\tilde{S}_{i}^{K}$ and $\tilde{S}_{j}^{K}$. Then
$\cos(\tilde{\alpha}_{ij}^{K})=-\frac{(\tilde{\boldsymbol{q}}_{i}^{K})^{T}\tilde{\boldsymbol{q}}_{j}^{K}}{\|\tilde{\boldsymbol{q}}_{i}^{K}\|\|\tilde{\boldsymbol{q}}_{j}^{K}\|}=-\frac{({\boldsymbol{q}}_{i}^{K})^{T}\mathbb{D}_{K}{\boldsymbol{q}}_{j}^{K}}{\|{\boldsymbol{q}}_{i}^{K}\|_{\mathbb{D}_{K}}\|{\boldsymbol{q}}_{j}^{K}\|_{\mathbb{D}_{K}}},\ \ i\neq j.$ (45)
Similarly, on the element $K$, we have the relation
$\cos(\alpha_{ij})=-{\boldsymbol{n}}_{i}\cdot{\boldsymbol{n}}_{j}=-\frac{{\boldsymbol{q}}_{i}\cdot{\boldsymbol{q}}_{j}}{\|{\boldsymbol{q}}_{i}\|\|{\boldsymbol{q}}_{j}\|}.$ (46)
It follows from (40) that
$\frac{h_{i}^{K}}{\sqrt{\rho_{max}(\mathbb{D}_{K}})}\leq\tilde{h}_{i}^{K}\leq\frac{h_{i}^{K}}{\sqrt{\rho_{min}(\mathbb{D}_{K}})},$ (47)
where $\rho_{min}(\mathbb{D}_{K})$ and $\rho_{max}(\mathbb{D}_{K})$ are the min and max eigenvalues of $\mathbb{D}_{K}$, respectively.
The difficulty with the finite element formulation mentioned in Section 2 is the coupling of the diffusion tensor with the flow field ${\boldsymbol{b}}$ in (27) as follows:
$(\mathbb{D}\nabla u_{h},\nabla({\boldsymbol{b}}\cdot\nabla\varphi^{h}))_{0,K}.$ (48)
For general ${\boldsymbol{b}}=(b_{1}(x_{1},x_{2}),b_{2}(x_{1},x_{2}))$, consider
$\frac{\partial{\boldsymbol{b}}}{\partial x_{i}}\otimes\mathbb{D}_{(i)},\ \ i=1,\dots,d,$ (49)
where $\mathbb{D}_{i}$ is the vector $i$th row of $\mathbb{D}$. An essential ingredient to the following analysis is that the tensor defined by
$\mathbb{K}=\sum_{i=1}^{d}\frac{\partial{\boldsymbol{b}}}{\partial x_{i}}\otimes\mathbb{D}_{(i)}$ (50)
is symmetric and positive definite. This highlights a significant relationship between $\frac{\partial{\boldsymbol{b}}}{\partial x_{i}}$ and the $i$th row of $\mathbb{D}$ when using SUPG. The added diffusion integral term in (6) can be written in general form as
$\tau_{K}\int_{K}(\mathbb{D}\nabla u)\cdot\nabla({\boldsymbol{b}}\cdot\nabla v)\textrm{d}{\boldsymbol{x}}$ $=\tau_{K}\sum_{i=1}^{d}\int_{K}(\nabla v)^{T}\left(\frac{\partial{\boldsymbol{b}}}{\partial x_{i}}\otimes\mathbb{D}_{(i)}\right)\nabla u\textrm{d}{\boldsymbol{x}}$ (51)
One may view $\mathbb{K}$ as a kind of weighted diffusion tensor.
The analysis of the numerical scheme relies on the following fundamental lemma [39].
Lemma 4.1.
If $\tilde{\boldsymbol{A}}$ in (32) is an $M$-matrix and has nonnegative row sums, then the numerical scheme satisfies DMP.
The following lemmas were used in [25] to analyze DMP with ${\boldsymbol{b}}={\boldsymbol{0}}$.
Lemma 4.2.
If the mesh satisfies the anisotropic nonobtuse angle condition (ANAC)
$0<\tilde{\alpha}_{ij}^{K}\leq\frac{\pi}{2},\ \ \forall i,j=1,\dots,d+1,i\neq j,\forall K\in\mathcal{T}_{h},$ (52)
then for any $\mathbb{F}_{K}$ of SPD-type, the matrix defined by
$\sum_{K\in\mathcal{T}_{h}}|K|(\nabla\varphi_{i})^{T}\mathbb{F}_{K}\nabla\varphi_{j},\ \ i=,1\dots N_{vi},j=1,\dots,N_{v}$ (53)
is an $M$-matrix and has nonnegative row sums. Moreover, for any element $K\in\mathcal{T}_{h}$, and $i,j=1,\dots,d+1$,
$(\nabla\varphi_{i})^{T}\mathbb{F}_{K}\nabla\varphi_{j}=\begin{cases}\displaystyle-\frac{\cos(\tilde{\alpha}_{ij}^{K})}{\tilde{h}_{i}^{K}\tilde{h}_{j}^{K}}\textrm{\ \ for \ \ }i\neq j\\ \displaystyle\frac{1}{(\tilde{h}_{i}^{K})^{2}}\textrm{\ \ for \ \ i = j}.\end{cases}$ (54)
When $d=2$, the case $i\neq j$ simplifies to
$|K|(\nabla\varphi_{i})^{T}\mathbb{F}_{K}\nabla\varphi_{j}=-\frac{\sqrt{\det(\mathbb{F}_{K}})}{2}\cot(\tilde{\alpha}_{ij}^{K}).$ (55)
We establish DMP in a series of steps. We first investigate spatial estimates, which are critical to understanding the required structure of the matrices ${\boldsymbol{A}}$ and ${\boldsymbol{M}}$. Temporal estimates are then derived in order to apply Lemma 4.1.
4.1 General flow forms
In this section we investigate the structure of $\mathbb{K}$ specified by ${\boldsymbol{b}}$. For $d=2$, we have
$\frac{\partial{\boldsymbol{b}}}{\partial x_{1}}\otimes\mathbb{D}_{(1)}=\begin{pmatrix}\displaystyle\frac{\partial b_{1}}{\partial x_{1}}D_{11}&\displaystyle\frac{\partial b_{1}}{\partial x_{1}}D_{12}\\ \displaystyle\frac{\partial b_{2}}{\partial x_{1}}D_{11}&\displaystyle\frac{\partial b_{2}}{\partial x_{1}}D_{12}\end{pmatrix}$ (56)
and
$\frac{\partial{\boldsymbol{b}}}{\partial x_{2}}\otimes\mathbb{D}_{(2)}=\begin{pmatrix}\displaystyle\frac{\partial b_{1}}{\partial x_{2}}D_{21}&\displaystyle\frac{\partial b_{1}}{\partial x_{2}}D_{22}\\ \displaystyle\frac{\partial b_{2}}{\partial x_{2}}D_{21}&\displaystyle\frac{\partial b_{2}}{\partial x_{2}}D_{22}\end{pmatrix}.$ (57)
There are two ways in which we consider $\mathbb{K}$. If (56) and (57) are of SPD-type, then $\mathbb{K}$ will be as well. If not, then we need to analyze $\mathbb{K}$ in its general form given by
$\mathbb{K}=\begin{pmatrix}\frac{\partial b_{1}}{\partial x_{1}}D_{11}+\frac{\partial b_{1}}{\partial x_{2}}D_{21}&\frac{\partial b_{1}}{\partial x_{1}}D_{12}+\frac{\partial b_{1}}{\partial x_{2}}D_{22}\\ \frac{\partial b_{2}}{\partial x_{1}}D_{11}+\frac{\partial b_{2}}{\partial x_{2}}D_{21}&\frac{\partial b_{2}}{\partial x_{1}}D_{12}+\frac{\partial b_{2}}{\partial x_{2}}D_{22}\end{pmatrix}.$ (58)
Hence, using the symmetric condition, we would require
$\frac{\partial}{\partial x_{1}}(b_{2}D_{11}-b_{1}D_{12})=\frac{\partial}{\partial x_{2}}(b_{1}D_{22}-b_{2}D_{21}).$ (59)
Assume first that these matrices are of SPD-type so that general form of $\mathbb{K}$ is of SPD-type.
Theorem 4.3.
Suppose $\mathbb{D}$, equations (56), and (57) are of SPD-type with constant real-valued entries and $D_{12}=D_{21}\neq 0$. Then there are functions $f_{1}(x_{2},t)$, $g_{1}(x_{2},t)$, $f_{2}(x_{1},t)$, and $g_{2}(x_{1},t)$ such that
$b_{1}(x_{1},x_{2},t)=\frac{D_{11}D_{22}f_{1}(x_{1},t)-D_{12}D_{21}f_{2}(x_{2},t)}{\det(\mathbb{D})}$ (60)
and
$b_{2}(x_{1},x_{2},t)$ $=\frac{D_{12}D_{22}(f_{1}(x_{1},t)-f_{2}(x_{2},t))}{\det(\mathbb{D})}.$ (61)
for $D_{22}\neq 0$. Likewise,
$b_{1}(x_{1},x_{2},t)=\frac{D_{11}D_{21}(g_{2}(x_{2},t)-g_{1}(x_{1},t))}{\det(\mathbb{D})}$ (62)
and
$b_{2}(x_{1},x_{2},t)=\frac{D_{11}D_{22}g_{2}(x_{2},t)-D_{12}D_{21}g_{1}(x_{1},t)}{\det(\mathbb{D})},$ (63)
whenever $D_{11}\neq 0$.
Proof.
Assume $D_{12}=D_{21}\neq 0$ and $D_{22}\neq 0$. The other case is very similar. Assuming symmetry as mentioned, we have
$\frac{\partial b_{1}}{\partial x_{1}}$ $=\left(\frac{D_{11}}{D_{12}}\right)\frac{\partial b_{2}}{\partial x_{1}}$ (64)
$b_{1}(x_{1},x_{2},t)$ $=\left(\frac{D_{11}}{D_{12}}\right)b_{2}(x_{1},x_{2},t)+f_{2}(x_{2},t),$ (65)
for some $f_{2}(x_{2},t)$. Additionally,
$\frac{\partial b_{1}}{\partial x_{2}}$ $=\left(\frac{D_{21}}{D_{22}}\right)\frac{\partial b_{2}}{\partial x_{2}}$ (66)
$b_{1}(x_{1},x_{2},t)$ $=\left(\frac{D_{21}}{D_{22}}\right)b_{2}(x_{1},x_{2},t)+f_{1}(x_{1},t),$ (67)
for some function $f_{1}(x_{1},t)$. Equating the two equations for $b_{1}$ gives
$\left(\frac{D_{11}}{D_{12}}\right)b_{2}(x_{1},x_{2},t)+f_{2}(x_{2},t)$ $=\left(\frac{D_{21}}{D_{22}}\right)b_{2}(x_{1},x_{2},t)+f_{1}(x_{1},t)$ (68)
$\left(\frac{D_{11}}{D_{12}}-\frac{D_{21}}{D_{22}}\right)b_{2}(x_{1},x_{2},t)$ $=f_{1}(x_{2},t)-f_{2}(x_{1},t)$ (69)
$b_{2}(x_{1},x_{2},t)$ $=\frac{D_{12}D_{22}(f_{1}(x_{1},t)-f_{2}(x_{2},t))}{\det(\mathbb{D})}.$ (70)
Plugging this into the equation for $b_{1}$, we have
$b_{1}(x_{1},x_{2},t)$ $=\left(\frac{D_{11}}{D_{12}}\right)\frac{D_{12}D_{22}(f_{1}(x_{1},t)-f_{2}(x_{2},t))}{\det(\mathbb{D})}+f_{2}(x_{2},t)$ (71)
$=\frac{D_{11}D_{22}(f_{1}(x_{1},t)-f_{2}(x_{2},t))+f_{2}(x_{2},t)(D_{11}D_{22}-D_{12}D_{21})}{\det(\mathbb{D})}$ (72)
$=\frac{D_{11}D_{22}f_{1}(x_{1},t)-D_{12}D_{21}f_{2}(x_{2},t)}{\det(\mathbb{D})}.$ (73)
Note that since $\mathbb{D}$ is of SPD-type, $\det(\mathbb{D})\neq 0$. ∎
      Remark 4.1.
If $D_{12}=D_{21}$ is not zero, the general form of ${\boldsymbol{b}}$ is given by
$\frac{1}{\det(\mathbb{D})}\left(D_{22}f_{1}(x_{1},t)\mathbb{D}_{(1)}-D_{12}f_{2}(x_{2},t)\mathbb{D}_{(2)}\right)$ (74)
for some functions $f_{1}$ and $f_{2}$ whenever $D_{22}\neq 0$. Similarly, if $D_{11}\neq 0$, then the general form of ${\boldsymbol{b}}$ is given by
$\frac{1}{\det(\mathbb{D})}\left(D_{11}g_{2}(x_{2},t)\mathbb{D}_{(2)}-D_{21}g_{1}(x_{1},t)\mathbb{D}_{(1)}\right)$ (75)
If $D_{11}$ and $D_{22}$ are both not zero, then $g_{1}=-\displaystyle\frac{D_{22}}{D_{21}}f_{1}$ and $g_{2}=-\displaystyle\frac{D_{12}}{D_{11}}f_{2}$.
Corollary 4.3.1.
If $\mathbb{D}=\textrm{diag}(D_{11},D_{22})$, with $D_{11}$ and $D_{22}\neq 0$, then there are functions $f$ and $g$ such that
${\boldsymbol{b}}(x_{1},x_{2},t)=\begin{pmatrix}\displaystyle\frac{f(x_{1},t)}{D_{11}}\\ \displaystyle\frac{g(x_{2},t)}{D_{22}}\end{pmatrix}$ (76)
and
$\mathbb{K}=\begin{pmatrix}\displaystyle\frac{\partial f}{\partial x_{1}}&0\\ 0&\displaystyle\frac{\partial g}{\partial x_{2}}\end{pmatrix}.$ (77)
Proof.
The form of ${\boldsymbol{b}}$ as shown in (76) follows from a similar argument as shown in the proof of the previous theorem. ∎
Based on the Remark 4.1, we require
$\frac{D_{22}}{\det(\mathbb{D})}\left(D_{11}\frac{\partial f_{1}}{\partial x_{1}}-D_{12}\frac{\partial f_{2}}{\partial x_{2}}\right).$ (78)
to share the same sign as the divergence of ${\boldsymbol{b}}$.
4.2 Mass and stiffness matrix structure: estimates on $\tau_{K}$
We first establish non-negative diagonal entries for the mass and stiffness matrices.
Lemma 4.4.
Suppose $\mathbb{K}_{K}$ and $\mathbb{D}_{K}$ are of SPD-type and $\mathbb{D}_{K}$ satisfies ANAC. If ${\boldsymbol{b}}$ has nonpositive divergence, then $A_{ii}$ and $M_{ii}$ are both strictly positive. If ${\boldsymbol{b}}$ has strictly positive divergence, then $A_{ii}>0$ if
$\tau_{K}>-\frac{\varphi_{i}}{{\boldsymbol{b}}\cdot\nabla\varphi_{i}}$ (79)
and $M_{ii}>0$ if
$\tau_{K}<\frac{4|K|}{(d+1)(d+2)}\left(\int_{K}(\nabla\cdot{\boldsymbol{b}})\varphi_{i}^{2}\textrm{d}{\boldsymbol{x}}\right)^{-1}.$ (80)
Proof.
We consider a patch of elements $\Omega_{i}$ that share a vertex ${\boldsymbol{a}}_{i}$. Recall the definition of $A_{ii}$ we have
$A_{ii}$ $=\sum_{K\in\Omega_{i}}\int_{K}\varphi_{i}{\boldsymbol{b}}\cdot\nabla\varphi_{i}+\tau_{K}({\boldsymbol{b}\cdot\nabla\varphi_{i}})^{2}+\tau_{K}\left(\varepsilon\mathbb{D}_{K}\nabla\varphi_{i}\right)\cdot\nabla({\boldsymbol{b}\cdot\nabla\varphi_{i})}\textrm{d}{\boldsymbol{x}}+|K|(\nabla\varphi_{i})^{T}\varepsilon\mathbb{D}_{K}\nabla\varphi_{i}$ (81)
We first concentrate on the first two integrals and make use of $\tau_{K}$ in order to bound these integrals below by zero. Note that by (43) the first integral is nonnegative if ${\boldsymbol{b}}$ has nonpositive divergence. Then
$\int_{K}(\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})+\tau_{K}({\boldsymbol{b}}\cdot\nabla\varphi_{i})^{2})\textrm{d}{\boldsymbol{x}}\geq 0.$ (82)
If ${\boldsymbol{b}}$ has strictly positive divergence, then (82) holds under the condition (79). Indeed, (43) implies
$\int_{K}\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}<0$ (83)
and hence ${\boldsymbol{b}}\cdot\nabla\varphi_{i}<0$ since $\varphi_{i}>0$ on element $K\in\Omega_{i}$ and $\tau_{K}>0$ for all $K\in\Omega$. By factoring out the convection term, the lower bound on $\tau_{K}$ given in equation (79) gives (82).
We now turn our attention to the third term in (81). Let $\lambda_{j}$ and ${\boldsymbol{\omega}}_{j}$ be the quadrature weights and nodes, respectively on element $K\in\Omega_{i}$. Then
$\int_{K}(\mathbb{D}\nabla\varphi_{i})\cdot\nabla({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}$ $\approx|K|\sum_{i=1}^{m}\lambda_{j}(\nabla({\boldsymbol{b}}({\boldsymbol{\omega}}_{j})\cdot\nabla\varphi_{i})^{T}\mathbb{D}({\boldsymbol{\omega}}_{j})\nabla\varphi_{i}$ (84)
$=|K|(\nabla\varphi_{i})^{T}\mathbb{K}_{K}(\nabla\varphi_{i})$ (85)
It follows from Lemma 4.2 that $A_{ii}>0$.
On the other hand,
$M_{ii}$ $=\sum_{K\in\Omega_{i}}(\varphi_{i},\varphi_{i}+\tau_{K}({\boldsymbol{b}}\cdot\nabla\varphi_{i}))_{0,K}$ (86)
$=\sum_{K\in\Omega_{i}}\int_{K}\left(\varphi_{i}^{2}+\tau_{K}\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\right)\textrm{d}{\boldsymbol{x}}$ (87)
$=\sum_{K\in\Omega_{i}}\left(\frac{2|K|}{(d+1)(d+2)}+\tau_{K}\int_{K}\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}\right).$ (88)
As we saw previously, the integral is nonnegative if ${\boldsymbol{b}}$ has nonpositive divergence. Otherwise, we apply the upper bound on $\tau_{K}$ given by (80). Hence, $M_{ii}>0$. ∎
      Remark 4.2.
To summarize the previous lemma, under the ANAC and SPD-type conditions, if ${\boldsymbol{b}}$ is nonpositive, then $A_{ii}$ and $M_{ii}>0$. Otherwise, we require
$-\frac{\varphi_{i}}{{\boldsymbol{b}\cdot\nabla\varphi_{i}}}<\tau_{K}<\frac{4|K|}{(d+1)(d+2)}\left(\int_{K}(\nabla\cdot{\boldsymbol{b}})\varphi_{i}^{2}\textrm{d}{\boldsymbol{x}}\right)^{-1}.$ (89)
It is important to keep in mind that this estimate is only for those elements associated with $A_{ii}$ on the interior of $\Omega$. This estimate highlights the significant relationship between $\tau_{K}$ and the convection term.
Lemma 4.5.
If $\mathbb{K}_{K}$ and ${\mathbb{D}}_{K}$ are of SPD-type, satisfy ANAC, and if the following two conditions are met
$\frac{|K|\|{\boldsymbol{b}}\|_{\infty,K}^{2}}{\rho}\leq\cos(\check{\alpha}_{ij})$ (90)
and
$\frac{2h_{i}^{K}\|{\boldsymbol{b}}\|_{\infty,K}}{\rho\sqrt{(d+1)(d+2)}}\leq\cos(\tilde{\alpha}_{ij}),$ (91)
where $\rho=\min\{\rho_{min}(\mathbb{K}_{K}),\rho_{min}(\varepsilon\mathbb{D}_{K})\}$, then $A_{ij}\leq 0$, for $i\neq j$.
Proof.
For the ${\boldsymbol{A}}$ matrix, we have
$A_{ij}=\sum_{K\in\Omega_{i}\cap\Omega_{j}}\int_{K}\varphi_{i}{\boldsymbol{b}}\cdot\nabla\varphi_{j}+\tau_{K}({\boldsymbol{b}\cdot\nabla\varphi_{j}})({\boldsymbol{b}\cdot\nabla\varphi_{i}})$ $+\tau_{K}\left(\varepsilon\mathbb{D}_{K}\nabla\varphi_{j}\right)\cdot\nabla({\boldsymbol{b}\cdot\nabla\varphi_{i})}\textrm{d}{\boldsymbol{x}}$ (92)
$+|K|(\nabla\varphi_{i})^{T}\varepsilon\mathbb{D}_{K}\nabla\varphi_{j}$ (93)
We investigate the three terms in the integral of $A_{ij}$. The first integral can be simplified using Cauchy-Schwarz and Holder’s inequalities along with (42),
$\int_{K}\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{j})\textrm{d}{\boldsymbol{x}}$ $\leq\sqrt{\frac{|K|}{(d+1)(d+2)}}\|{\boldsymbol{b}}\|_{\infty,K}\|\nabla\varphi_{i}\|_{2,K}$ (94)
$=\sqrt{\frac{2}{(d+1)(d+2)}}\left(\frac{|K|\|{\boldsymbol{b}}\|_{\infty,K}}{h_{i}^{K}}\right).$ (95)
For the second term, we again use Cauchy-Schwarz and Holder’s inequality and have the following bound on the convection integral
$\int_{K}({\boldsymbol{b}}\cdot\nabla\varphi_{i})({\boldsymbol{b}}\cdot\nabla\varphi_{j})\textrm{d}{\boldsymbol{x}}$ $\leq\|{\boldsymbol{b}}\|_{\infty,K}^{2}\frac{|K|}{h_{i}^{K}h_{j}^{K}}.$ (96)
We have
$A_{ij}$ $\leq\sum_{K\in\Omega_{i}\cap\Omega_{j}}\left(\frac{\sqrt{2}|K|}{h_{j}^{K}\sqrt{(d+1)(d+2)}}\|{\boldsymbol{b}}\|_{\infty,K}+\tau_{K}\|{\boldsymbol{b}}\|_{\infty,K}^{2}\frac{|K|}{h_{i}^{K}h_{j}^{K}}-\tau_{K}\frac{\cos(\check{\alpha}_{ij}^{K})}{\check{h}_{i}^{K}\check{h}_{j}^{K}}-|K|\frac{\cos(\tilde{\alpha}_{ij}^{K})}{\tilde{h}_{i}^{K}\tilde{h}_{j}^{K}}\right)$ (97)
$\leq\sum_{K\in\Omega_{i}\cap\Omega_{j}}\frac{|K|\|{\boldsymbol{b}}\|_{\infty,K}}{h_{j}^{K}}\left(\frac{\sqrt{2}}{\sqrt{(d+1)(d+2)}}+\frac{\tau_{K}\|{\boldsymbol{b}}\|_{\infty,K}}{h_{i}^{K}}\right)$ (98)
$\hskip 144.54pt-\frac{1}{h_{i}^{K}h_{j}^{K}}\left(\tau_{K}\cos(\check{\alpha}_{ij}^{K})\rho_{min}(\mathbb{K}_{K})+|K|\cos(\tilde{\alpha}_{ij}^{K})\rho_{min}(\varepsilon\mathbb{D}_{K})\right)$ (99)
$\leq\sum_{K\in\Omega_{i}\cap\Omega_{j}}\frac{|K|\|{\boldsymbol{b}}\|_{\infty,K}}{h_{j}^{K}}\left(\frac{\sqrt{2}}{\sqrt{(d+1)(d+2)}}+\frac{\tau_{K}\|{\boldsymbol{b}}\|_{\infty,K}}{h_{i}^{K}}\right)$ (100)
$\hskip 144.54pt-\frac{\min\{\rho_{min}(\mathbb{K}_{K}),\rho_{min}(\varepsilon\mathbb{D}_{K})\}}{h_{i}^{K}h_{j}^{K}}\left(\tau_{K}\cos(\check{\alpha}_{ij})+|K|\cos(\tilde{\alpha}_{ij})\right)$ (101)
By (90) and (91), $A_{ij}\leq 0$. ∎
Lemma 4.6.
If
$\int_{K}\varphi_{j}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}\geq 0$ (102)
or
$\tau_{K}\leq-\frac{|K|}{\displaystyle(d+1)(d+2)}\left(\int_{K}\varphi_{j}(\boldsymbol{b}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}\right)^{-1},$ (103)
then $M_{ij}\geq 0$.
Proof.
The is immediate based on the definition of $M_{ij}$ and (21). ∎
Combining the previous two lemmas, we have the following result.
Lemma 4.7.
Let $\mathbb{K}_{K}$ and $\mathbb{D}_{K}$ be of SPD-type, satisfy ANAC, and (90) and (91) be satisfied. The following nonnegative condition
$M_{ij}-\Delta t(1-\theta)A_{ij}\geq 0$ (104)
holds for the off-diagonal entries if the convection integral is nonnegative or $\tau_{K}$ satisfies (103).
4.3 Temporal estimates
In this section we derive upper and lower bounds on the fixed time step $\Delta t$ so that DMP is satisfied.
Lemma 4.8.
Suppose $\Delta t$ satisfies
$\Delta t\geq-\frac{1}{\theta\sqrt{(d+1)(d+2)}}\max_{K\in\mathcal{T}_{h}}\max_{\underset{i\neq j}{i,j=1,\dots,d+1}}\left(\frac{|K|h_{j}^{K}\left(h_{i}^{K}+2\tau_{K}\sqrt{(d+1)(d+2)}\|{\boldsymbol{b}}\|_{\infty,K}\right)}{2|K|h_{i}^{K}\|{\boldsymbol{b}}\|_{\infty,K}+\tau_{K}h_{i}^{K}\|{\boldsymbol{b}}\|_{\infty,K}^{2}-\rho(\tau_{K}\cos(\check{\alpha}_{ij})+|K|\cos(\tilde{\alpha}_{ij}))}\right),$ (105)
where $\rho=\min\{\rho_{min}(\mathbb{K}_{K}),\rho_{min}(\varepsilon\mathbb{D}_{K})\}$. Suppose also that $\mathbb{K}_{K}$ and $\mathbb{D}_{K}$ are of SPD-type, satisfy ANAC, and satisfy (90) and (91) (without equality). If, for each element $K$, ${\boldsymbol{b}}$ has nonpositive divergence on or $\tau_{K}$ satisfies (89), then $\tilde{\boldsymbol{A}}$ is an $M$-matrix.
Proof.
From Lemma 4.4, the diagonal entries of ${\boldsymbol{M}}+\theta\Delta t{\boldsymbol{A}}$ is strictly positive. We show
$M_{ij}+\theta\Delta tA_{ij}\leq 0,\ \ \forall i\neq j,i=1,\dots,N_{vi},j=1,\dots,N_{v},$ (106)
since this is equivalent to showing that $M+\theta\Delta tA$ is a $Z$-matrix and an $M$-matrix is a $Z$-matrix that is inverse positive.
For the ${\boldsymbol{M}}$ matrix, we have
$M_{ij}$ $=\sum_{K\in\Omega_{i}\cap\Omega_{j}}\left(\frac{|K|}{(d+1)(d+2)}+\tau_{K}\int_{K}\varphi_{j}({\boldsymbol{b}}\cdot\nabla\varphi_{i})\textrm{d}{\boldsymbol{x}}\right)$ (107)
The second integral term can be treated similar as (95). Now, we obtain the upper bound for $M_{ij}$,
$M_{ij}\leq\sum_{K\in\Omega_{i}\cap\Omega_{j}}\frac{|K|}{h_{i}^{K}[(d+1)(d+2)]^{3/2}}\left(h_{i}^{K}\sqrt{(d+1)(d+2)}+\tau_{K}\sqrt{2}(d+1)(d+2)\|{\boldsymbol{b}}\|_{\infty,K}\right).$ (108)
Recall we have the following bound on $A_{ij}$
$\begin{split}A_{ij}&\leq\sum_{K\in\Omega_{i}\cap\Omega_{j}}\frac{|K|\|{\boldsymbol{b}}\|_{\infty,K}}{h_{j}^{K}}\left(\frac{2}{\sqrt{(d+1)(d+2)}}+\frac{\tau_{K}\|{\boldsymbol{b}}\|_{\infty,K}}{h_{i}^{K}}\right)-\frac{\rho}{h_{i}^{K}h_{j}^{K}}\left(\tau_{K}\cos(\check{\alpha}_{ij})+|K|\cos(\tilde{\alpha}_{ij})\right)\\ &=\sum_{K\in\Omega_{i}\cap\Omega_{j}}\frac{1}{h_{i}^{K}h_{j}^{K}\sqrt{(d+1)(d+2)}}\left(2h_{i}^{K}h_{j}^{K}|K|\|{\boldsymbol{b}}\|_{\infty,K}+h_{j}^{K}|K|\tau_{K}\sqrt{(d+1)(d+2)}\|{\boldsymbol{b}}\|_{\infty,K}^{2}\right.\\ &\hskip 198.7425pt\left.-\rho\sqrt{(d+1)(d+2)}\left(\tau_{K}\cos(\check{\alpha}_{ij})+|K|\cos(\tilde{\alpha}_{ij})\right)\right)\end{split}$ (109)
Thus, $\tilde{\boldsymbol{A}}$ in (33) is a $Z$-matrix using the assumption on $\Delta t$.
By (33), we can write
$\tilde{\boldsymbol{A}}=\begin{bmatrix}M_{11}+\theta\Delta tA_{11}&M_{12}+\theta\Delta tA_{12}\\ 0&I_{N_{v}-N_{vi}}\end{bmatrix}.$
Since $\tilde{\boldsymbol{A}}$ is a $Z$-matrix, the block $M_{11}+\theta\Delta tA_{11}$ is also a $Z$-matrix. First, we show $M_{11}$ is positive-definite. Let ${\boldsymbol{v}}=(v_{1},v_{2},\dots,v_{N_{vi}})$ be a nonzero vector in $\mathbb{R}^{N_{vi}}$ and write $v^{h}=\sum_{j=1}^{N_{vi}}v_{j}\varphi_{j}\in U_{h}^{0}$. Using the assumption on the divergence of ${\boldsymbol{b}}$ being either nonnegative of $\tau_{K}$ satisfying (89), ${\boldsymbol{v}}^{T}M_{11}{\boldsymbol{v}}>0$ and ${\boldsymbol{v}}^{T}A_{11}{\boldsymbol{v}}>0$ (this is a similar calculation as shown in Lemma 4.4). Therefore, $M_{11}+\theta\Delta tA_{11}$ is an $M$-matrix. We have
$\tilde{\boldsymbol{A}}^{-1}=\begin{bmatrix}(M_{11}+\theta\Delta A_{11})^{-1}&-(M_{11}+\theta\Delta tA_{11})^{-1}(M_{12}+\theta\Delta tA_{12})\\ 0_{N_{v}}&I_{N_{v}-N_{vi}}\end{bmatrix},$ (110)
which implies $\tilde{\boldsymbol{A}}$ is an $M$-matrix since $\tilde{\boldsymbol{A}}^{-1}\geq 0$ (note that $M_{12}+\theta\Delta tA_{12}\leq 0$). ∎
Lemma 4.9.
(104) holds along the diagonal entries if the mesh generated by $\mathbb{D}_{K}$ and $\mathbb{K}_{K}$ satisfy ANAC and are of SPD-type, and $\Delta t$ satisfies the bound
$\Delta t\leq\frac{1}{(1-\theta)(d+1)(d+2)}\min_{K\in\mathcal{T}_{h}}\min_{i,j=1,\dots,d+1}\frac{(\tilde{h}_{i}^{K}\check{h}_{i}^{K})^{2}(\tau_{K}(d+1)(d+2)\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})+2)}{|K|(\check{h}_{i}^{K})^{2}+(\tilde{h}_{i}^{K}\check{h}_{i}^{K})^{2}(\varphi_{i}({\boldsymbol{b}}\cdot\nabla\varphi_{i})+\tau_{K}({\boldsymbol{b}}\cdot\nabla\varphi_{i})^{2})+\tau_{K}(\tilde{h}_{i}^{K})^{2}}$ (111)
Proof.
From (88) and (81), along with (54), it’s clear that if (111) holds, then $M_{ii}-(1-\theta)A_{ii}\geq 0$. ∎
4.4 Metric tensor for DMP
We now develop the metric tensor to control the mesh so that the numerical solution satisfies DMP. Assume the temporal estimates from the previous section. As mentioned in the proof of Theorem 3.1 of [26], $\tilde{\boldsymbol{A}}^{-1}\geq 0$ and $\tilde{\boldsymbol{A}}$ is an $M$-matrix. To see that $\tilde{\boldsymbol{A}}$ has nonnegative row sums, note that $\sum_{j}\varphi_{j}=1$. Therefore, the solution to (32) satisfies
$\underset{i=1,\dots,(N+1)N_{v}}{\max}(u_{h})_{i}=\max\left\{0,\underset{i\in S({\boldsymbol{f}}^{+})}{\max}(u_{h})_{i}\right\},$ (112)
where $S({\boldsymbol{f}}^{+})$ is the set of indices where $f_{i}$ is nonnegative. Piecewise linear functions attain their maximum value at the vertices of element $K$. Hence, when $f({\boldsymbol{x}},t)\leq 0$, we have
$\underset{n=0,\dots,N}{\max}\underset{{\boldsymbol{x}}\in\overline{\Omega}}{\max}\ u_{h}^{n}({\boldsymbol{x}})=\max\left\{0,\underset{n=0,\dots,N}{\max}\underset{{\boldsymbol{x}}\in\overline{\Omega}}{\max}\ u_{h}^{n}({\boldsymbol{x}}),\underset{{\boldsymbol{x}\in\overline{\Omega}}}{\max}\ u_{h}^{0}({\boldsymbol{x}})\right\},$ (113)
where
$u_{h}^{n}=\sum_{j=1}^{N_{vi}}u_{j}^{n}\varphi_{j}({\boldsymbol{x}})+\sum_{j=N_{vi}+1}^{N_{v}}u_{j}^{n}\varphi_{j}({\boldsymbol{x}}),\ \ n=0,\dots,N.$ (114)
Theorem 4.10.
Suppose the two meshes generated by $\mathbb{D}_{K}$ and $\mathbb{K}_{K}$ satisfy ANAC, are of SPD-type, and satisfy the step size bounds (105) and (111). If ${\boldsymbol{b}}$ has nonpositive divergence or $\tau_{K}$ satisfies (89), then the SUPG scheme satisfies DMP.
Since two different metric tensors will be used on any given node, we follow [40] and taken into consideration for every time step, we use metric intersection in order to generate the anisotropic mesh.
Figure 11: Geometric illustration of intersecting ellipses $\mathcal{E}$ and $\mathcal{E}^{\prime}$, resulting in $\mathcal{E}\cap\mathcal{E}^{\prime}$.
This is equivalent to finding the largest ellipsoid included in the intersection of the two ellipses $\mathcal{E}$ and $\mathcal{E}^{\prime}$, the geometric representations of symmetric, positive definite $\mathcal{M}$ and $\mathcal{M}^{\prime}$, respectively (see Figure 11). By our assumptions on $\mathcal{M}$ and $\mathcal{M}^{\prime}$, $\mathcal{N}=\mathcal{M}^{-1}\mathcal{M}^{\prime}$ is diagonalizable by the similarity transform. Denote the shared normalized eigenvectors of $\mathcal{M}$ and $\mathcal{M}^{\prime}$ by ${\boldsymbol{\lambda}}_{1}$ and ${\boldsymbol{\lambda}}_{2}$ and write
$\mathcal{M}_{int}=\mathcal{M}\cap\mathcal{M}^{\prime}=\Lambda^{-T}\textrm{diag}\left(\max\{\alpha_{1},\beta_{1}\},\max\{\alpha_{2},\beta_{2}\}\right)\Lambda^{-1}.$ (115)
Figure 12 shows the coordinate transformation of a typical element $K$ to $\tilde{K}$ and $\check{K}$.
[…]
Figure 12: Coordinate transformation of a typical element $K$ to $\tilde{K}$ and $\check{K}$.
4.5 Numerical Results
In this section we present our numerical findings for using different metric tensors. We start with the metric tensors $\mathcal{M}_{DMP}=\theta\mathbb{D}^{-1}$ and $\mathcal{M}_{\mathbb{K}}=\theta^{\prime}\mathbb{K}^{-1}$ and define the metric tensor $\mathcal{M}_{int}=\mathcal{M}_{DMP}\cap\mathcal{M}_{\mathbb{K}}$. We use the implicit Euler method in time for all computations (that is, $\theta=1$) and compare the results with the classic Galerkin method with SUPG using fixed mesh (FM-SUPG).
      Example 4.1.
Consider the manufactured solution on the domain $\Omega_{T}=[0,1]^{2}\times[0.2,0.3]$ to (1) given by
$u(x,y,t)=(1+\exp(C(x+y-t))^{-1},\ \ C>0$ (116)
with diffusion tensor defined by
$\mathbb{D}=\begin{pmatrix}50.5&49.5\\ 49.5&50.5\end{pmatrix},$ (117)
and time-dependent flow given by
${\boldsymbol{b}}(x,y,t)=(x+t,y+t).$ (118)
In the computations, we have used $\varepsilon=10^{-6}$ and $\Delta t=10^{-2}$ for this example. Based on the maximum principle, the solution should satisfy $\underset{\Omega_{T}}{\min}\ u=0$.
The diffusion tensor has eigenvalues $\lambda_{1}=100$ and $\lambda_{2}=1$, which indicates that the diffusion in horizontal direction is 100 times faster than that in the vertical direction. For the given flow in (118), the tensor $\mathbb{K}$ defined in (50) is the same as $\mathbb{D}$. Therefore, the metric tensor $\mathcal{M}_{DMP}$ is the same as $\mathcal{M}_{\mathbb{K}}$ for this example.
Table 5 presents the minimum values of the numerical solutions obtained from different meshes. As can be seen, the results from $\mathcal{M}_{int}$ and $\mathcal{M}_{DMP}$ are nearly identical. Even the results from the fixed mesh show satisfaction of DMP for the bottom shelf. However, the graph in Figure 13(a) shows significant artificial oscillation near the top shelf when using the fixed mesh approach. On the other hand, as shown in Figure 13(b), when using $\mathcal{M}_{int}$ or $\mathcal{M}_{DMP}$, the artificial oscillation at the top shelf is essentially eliminated.
Table 5: Example 4.1. $\underset{\Omega_{T}}{\min}\ u$ with $\Delta t=10^{-2}$.
$N$ $\mathcal{M}_{int}$ $\mathcal{M}_{DMP}$ Fixed mesh
16200 -4.4374e-05 -4.4373e-05 -2.8664e-07
24200 -3.1365e-06 -3.1367e-06 -2.2583e-08
34322 -1.5736e-07 -1.5735e-07 -1.1880e-09
Refer to caption (a): Fixed mesh           Refer to caption (b): $\mathcal{M}_{int}$
Figure 13: Example 4.1 – side view of solutions for a fixed mesh and $\mathcal{M}_{int}$ with $N=34322$, $\varepsilon=10^{-6}$, and $\Delta t=10^{-2}$.
      Example 4.2.
In this example, we consider the metric tensor $\mathcal{M}_{DMP+adapt}$ developed in [25] that is based on both DMP satisfaction as well as minimization of the $H^{1}$-seminorm of the interpolation error. Correspondingly, our intersection metric tensor is $\mathcal{M}_{int+adapt}=\mathcal{M}_{DMP+adap}\cap\mathcal{M}^{\prime}$ where $\mathcal{M}^{\prime}=\theta^{\prime}\mathbb{K}^{-1}$. The anisotropic diffusion tensor is given by
$\mathbb{D}=\begin{pmatrix}1&1\\ 1&3\end{pmatrix}$ (119)
and the flow is given by
${\boldsymbol{b}}(x,y,t)=(x+y+t+8,x+3y+t+16),$ (120)
for $\Omega_{T}=[0,1.5]^{2}\times[0,0.015]$. For the given ${\boldsymbol{b}}$ and $\mathbb{D}$, we have
$\mathbb{K}=\begin{pmatrix}2&4\\ 4&10\end{pmatrix}.$ (121)
We first use $N=20,402$ elements in the mesh and compute the numerical solutions for $\varepsilon=10^{-3}$ with time step size $\Delta t=10^{-4}$. Figure 14 shows the side view of the solution using fixed mesh, $\mathcal{M}_{DMP+adapt}$ mesh and $\mathcal{M}_{int+adapt}$ mesh, respectively. As can be seen, $\mathcal{M}_{DMP+adapt}$ mesh and $\mathcal{M}_{int+adapt}$ mesh provide comparable results and both perform much better than the fixed mesh SUPG approach.
Refer to caption (a): Fixed mesh Refer to caption (b): $\mathcal{M}_{DMP+adapt}$ Refer to caption (c): $\mathcal{M}_{int+adapt}$
Figure 14: Example 4.2 – side view with $\varepsilon=10^{-3}$ for different metric tensors and $\Delta t=10^{-4}$, $N=20,402$.
To see the advantage of $\mathcal{M}_{DMP+adapt}$ mesh over $\mathcal{M}_{int+adapt}$ mesh, we compute the numerical solutions at different $N$ values with larger time step $\Delta t=10^{-3}$. The minimum values of the numerical solutions are summarized in Table 6. As can be seen, $\mathcal{M}_{int+adapt}$ provides the best results.
Table 6: Example 4.2. Results for $\mathcal{M}_{int+adapt}$, $\mathcal{M}_{DMP+adapt}$, and fixed mesh for varying $N$ with $\varepsilon=10^{-3}$ and $\Delta t=10^{-3}$.
$N$ Fixed Mesh $\mathcal{M}_{DMP+adapt}$ $\mathcal{M}_{int+adapt}$
3362 -2.4050e-01 -2.3330e-03 -3.6297e-03
13122 -1.6762e-01 -1.3796e-03 -6.8634e-04
29282 -1.0677e-01 -9.3509e-04 -1.8547e-04
5 Summary and Conclusions
In this paper, we have developed a moving mesh SUPG (MM-SUPG) method that integrates the moving mesh partial differential equation (MMPDE) methods with streamline upwind Petrov-Galerkin (SUPG) method for convection-diffusion problems. MM-SUPG method takes the advantage of the SUPG method to stabilize the numerical solution and the moving mesh method to adapt the mesh to improve the efficiency of the computation and accuracy of the solution. We considered convection-dominated problems where sharp layers exist in the solution. Both time-dependent and time-independent flows are investigated. And the method is designed to work with both isotropic and anisotropic diffusion tensors.
For isotropic diffusion case, our MM-SUPG method shows a $O(h^{1})$ convergence rate in the $H^{1}$-seminorm and $O(h^{1.5})$ convergence in the $L^{2}$ norm (see Figures 1 and 4). Even with moderately large $N$, we see in Figure 5 that oscillations can still be present when using moving mesh with standard Galerkin finite element method (MM-FEM). However, when using MM-SUPG method, the oscillations are significantly reduced and the solution is much smoother. The results in Examples 3.1 and 3.2 demonstrate the advantage of MM-SUPG over fixed mesh with SUPG (FM-SUPG) and moving mesh without SUPG (MM-FEM).
For anisotropic diffusion case, we have developed conditions for mesh and time step size so that the numerical solution satisfies the discrete maximum principle. Based on the diffusivity tensor $\mathbb{D}$ and the flow vector $\mathbf{b}$, we have introduced a weighted diffusion tensor $\mathbb{K}$ in (50) and developed a new metric tensor, $\mathcal{M}_{int}$ in (115) to control the mesh. Numerical results in Examples 4.1 and 4.2 demonstrate that $\mathcal{M}_{int}$ mesh provides better solution than other meshes in terms of DMP satisfaction.
In summary, the MM-SUPG method takes the advantage of both SUPG method and moving mesh method to provide a stable and accurate solution for convection-diffusion problems. Together with conditions for mesh and time step size, the MM-SUPG method can provide a solution that satisfies the discrete maximum principle, especially when the diffusion tensor is anisotropic.